OpenStudy (anonymous):

URGENT! WILL GIVE MEDAL! PLEASE SHOW STEPS OF WORKING OUT! Find x if 3-x^2, -2x and 1-x are consecutive terms of an arithmetic sequence.

OpenStudy (academicgurusinc):

Use the concept of a generic arithmetic sequence; the difference between two sequential terms must be equal to d Therefore, 3-x^2+d=-2x and, -2x +d = 1-x Solve for d: 3-x^2+d=-2x ---> d=x^2 - 2x-3 -2x +d = 1-x ---> d= 1 +x Now solve for x: x^2 - 2x-3= 1 +x ---> x^2-3x-4=0 Factor out: (x-4)(x+1) Therefore x= 4, -1

OpenStudy (academicgurusinc):

We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)